5. Using the GC mentioned in question 4 and the values listed on the chromatogram for the height and width (at half height) o1400- 1200- 1000- A-11,40 800 - Response - Mill Volts 12.07 Height: 72 mm Width: 9 mm 800 - Height: 125 mm Width: 8 mm 400- HProton NMR Spectrum for Peak A on Gas Chromatogram: ЗН 1H 2H 1H 9 8 7 6 5 4 3 2 1 0 PPM Expansion of downfield signals: 1H 2HProton NMR Spectrum for Peak B on Gas Chromatogram: ЗН 2H 2H i o 8 5 PPM Downfield signals are both doublets.Proton NMR Spectrum for Peak Con Gas Chromatogram: 3H 2H 1H 1H 9 8 7 6 5 3 2 0 PPM Expansion of downfield signals: 2H 1H 1H t
Answer #1

Solution:

Total height of three peaks = 125 + 72 + 10 = 207 mm

Percent composition of A = (125/207) x 100% = 60.38%

Percent composition of B = (72/207)x 100% = 34.78%

Percent composition of C = (10/207)x 100% = 4.8 %

NMR analysis:

7.3 ppm Ha 2.5 ppm 7.5 ppm Hb. CH3 =0 7.5 ppm Hc Hd 7.9 ppm

Peak B is para nitro Toluene:

Para isomer has two doublets only due to two identical set of protons having J= 8 Hz. This spectra can used to decide the position of nitro group.

7.5 ppm Ha 8.2 ppm 2.3 ppm H3C. Hb 7.5 ppm Ha NO Hb 8.2 ppm

Peak C is meta nitro Toluene

2.3 ppm CH3 7.5 ppm 8.05 ppm Hd На 7.45 ppm Hc NO Hb 8.1 ppm